# StochCA: A Novel Approach for Exploiting Pretrained Models with Cross-Attention

### v1

@@ 4.4 Experimental Results
ERM
$86.9$ $79.2$ $73.0$ $\mathbb{47.8}$ $71.7$
${}^{\dagger}$
$83.9$ $79.5$ $72.3$ $27.1$ $65.7$
${}^{\dagger}$
$82.3$ $79.1$ $72.3$ $34.3$ $67.0$
${}^{\dagger}$
$82.6$ $79.6$ $71.9$ $35.0$ $67.3$
${}^{\dagger}$
$86.6$ $78.9$ $74.5$ $41.3$ $70.3$
${}^{\dagger}$
$86.2$ $79.2$ $74.5$ $47.7$ $71.9$
${}^{\dagger}$
$87.0$ $78.7$ $\mathbb{74.8}$ $47.6$ $72.0$
${}^{\dagger}$
$\mathbb{87.4}$ $79.1$ $74.5$ $46.7$ $72.0$
${}^{\dagger}$
$86.9$ $79.8$ $74.5$ $47.4$ $72.2$
${}^{\dagger}$ $88.1$ $80.4$ $76.0$ $48.3$ $73.2$
DoPrompt
$86.7$ $78.9$ $72.8$ $47.4$ $71.5$
StochCA
$\mathbb{87.4}$ $\mathbb{80.0}$ $74.6$ $47.7$ $\mathbb{72.4}$
@@
${}^{\ddagger}$
$83.4$ $75.9$ $66.4$ $44.4$ $67.5$
${}^{\ddagger}$
$95.6$ $82.2$ $82.5$ $54.0$ $78.6$
StochCA
$94.8$ $82.1$ $82.2$ $58.0$ $79.3$
MIRO+StochCA
$\mathbb{96.4}$ $\mathbb{83.0}$ $\mathbb{85.1}$ $\mathbb{60.0}$ $\mathbb{81.1}$
Table 2: Domain generalization performance using ImageNet and CLIP ViT backbones. The best performance for each dataset is highlighted in bold. Performances marked with ${}^{\dagger}$ and ${}^{\ddagger}$ refer to the results reported in [29] and [14], respectively.
@@
ERM
$90.3_{\pm 0.8}$ $84.2_{\pm 0.5}$ $99.3_{\pm 0.2}$ $73.9_{\pm 3.2}$ $86.9$
${}^{\dagger}$
$89.1_{\pm 0.7}$ $79.1_{\pm 1.1}$ $98.3_{\pm 0.1}$ $68.2_{\pm 5.1}$ $83.9$
${}^{\dagger}$
$87.3_{\pm 0.6}$ $80.1_{\pm 0.5}$ $98.5_{\pm 0.2}$ $63.1_{\pm 2.9}$ $82.3$
${}^{\dagger}$
$87.5_{\pm 0.4}$ $78.6_{\pm 0.9}$ $98.0_{\pm 0.2}$ $66.3_{\pm 1.7}$ $82.6$
${}^{\dagger}$
$90.1_{\pm 0.6}$ $80.5_{\pm 1.0}$ $98.8_{\pm 0.1}$ $77.1_{\pm 1.5}$ $86.6$
${}^{\dagger}$
$89.7_{\pm 0.3}$ $80.2_{\pm 1.0}$ $98.7_{\pm 0.2}$ $76.1_{\pm 1.6}$ $86.2$
${}^{\dagger}$
$90.4_{\pm 0.5}$ $83.0_{\pm 0.7}$ $98.7_{\pm 0.0}$ $76.0_{\pm 0.2}$ $87.0$
${}^{\dagger}$
$91.1_{\pm 0.1}$ $84.0_{\pm 0.3}$ $99.2_{\pm 0.2}$ $75.4_{\pm 0.5}$ $87.4$
${}^{\dagger}$
$89.9_{\pm 0.5}$ $80.8_{\pm 0.5}$ $98.7_{\pm 0.1}$ $77.1_{\pm 0.7}$ $86.6$
${}^{\dagger}$ $91.1_{\pm 0.3}$ $83.0_{\pm 0.2}$ $99.6_{\pm 0.0}$ $78.7_{\pm 0.8}$ $88.1$
DoPrompt
$89.9_{\pm 2.2}$ $84.1_{\pm 0.7}$ $98.9_{\pm 0.0}$ $74.1_{\pm 3.3}$ $86.7$
StochCA
$90.8_{\pm 0.6}$ $82.9_{\pm 0.7}$ $98.9_{\pm 0.1}$ $77.1_{\pm 2.4}$ $87.4$
@@
ERM
$96.2_{\pm 0.5}$ $64.4_{\pm 0.9}$ $77.3_{\pm 0.3}$ $79.0_{\pm 1.0}$ $79.2$
${}^{\dagger}$
$96.2_{\pm 0.5}$ $65.7_{\pm 0.8}$ $76.8_{\pm 0.4}$ $77.7_{\pm 0.9}$ $79.5$
${}^{\dagger}$
$96.3_{\pm 0.2}$ $66.3_{\pm 0.5}$ $76.5_{\pm 0.6}$ $77.3_{\pm 0.5}$ $79.1$
${}^{\dagger}$
$97.7_{\pm 0.3}$ $66.5_{\pm 0.5}$ $76.0_{\pm 1.0}$ $78.0_{\pm 0.7}$ $79.6$
${}^{\dagger}$
$97.1_{\pm 0.5}$ $65.2_{\pm 0.7}$ $75.1_{\pm 1.0}$ $78.3_{\pm 1.0}$ $78.9$
${}^{\dagger}$
$96.2_{\pm 0.7}$ $66.0_{\pm 0.3}$ $76.8_{\pm 0.6}$ $78.0_{\pm 0.9}$ $79.2$
${}^{\dagger}$
$95.7_{\pm 0.5}$ $65.9_{\pm 0.3}$ $76.3_{\pm 0.8}$ $78.3_{\pm 1.0}$ $78.9$
${}^{\dagger}$
$97.1_{\pm 0.9}$ $65.4_{\pm 0.2}$ $76.1_{\pm 0.4}$ $78.0_{\pm 0.3}$ $79.1$
${}^{\dagger}$
$96.8_{\pm 0.1}$ $65.6_{\pm 1.0}$ $77.3_{\pm 0.6}$ $79.3_{\pm 1.5}$ $79.8$
${}^{\dagger}$ $97.2_{\pm 0.1}$ $67.4_{\pm 0.5}$ $77.4_{\pm 1.0}$ $79.7_{\pm 0.5}$ $80.4$
DoPrompt
$96.2_{\pm 0.6}$ $64.5_{\pm 0.8}$ $76.8_{\pm 1.5}$ $78.0_{\pm 1.1}$ $78.9$
StochCA
$96.3_{\pm 0.3}$ $65.9_{\pm 0.8}$ $77.9_{\pm 0.6}$ $79.9_{\pm 0.7}$ $80.0$
Table 3: Detailed results on PACS (left) and VLCS (right). Performances marked with ${}^{\dagger}$ refer to the results reported in [29].
@@
ERM
$69.8_{\pm 0.5}$ $59.4_{\pm 1.0}$ $79.6_{\pm 0.5}$ $83.1_{\pm 0.4}$ $73.0$
${}^{\dagger}$
$69.5_{\pm 0.1}$ $57.1_{\pm 0.3}$ $79.1_{\pm 0.1}$ $83.3_{\pm 0.1}$ $72.3$
${}^{\dagger}$
$68.2_{\pm 0.4}$ $56.8_{\pm 0.4}$ $79.8_{\pm 0.4}$ $84.2_{\pm 0.0}$ $72.3$
${}^{\dagger}$
$68.7_{\pm 0.9}$ $55.8_{\pm 0.4}$ $79.2_{\pm 0.0}$ $83.9_{\pm 0.2}$ $71.9$
${}^{\dagger}$
$71.2_{\pm 0.7}$ $60.9_{\pm 0.4}$ $81.1_{\pm 0.3}$ $84.7_{\pm 0.3}$ $74.5$
${}^{\dagger}$
$72.3_{\pm 0.7}$ $60.9_{\pm 1.0}$ $80.4_{\pm 0.4}$ $84.3_{\pm 0.0}$ $74.5$
${}^{\dagger}$
$72.3_{\pm 0.0}$ $60.0_{\pm 1.1}$ $82.2_{\pm 0.2}$ $84.7_{\pm 0.1}$ $74.8$
${}^{\dagger}$
$72.7_{\pm 0.4}$ $59.7_{\pm 0.5}$ $80.9_{\pm 0.3}$ $84.6_{\pm 0.2}$ $74.5$
${}^{\dagger}$
$71.8_{\pm 0.2}$ $60.7_{\pm 1.0}$ $81.4_{\pm 0.3}$ $84.3_{\pm 0.1}$ $74.5$
${}^{\dagger}$ $72.7_{\pm 0.5}$ $62.3_{\pm 0.6}$ $83.2_{\pm 0.1}$ $85.9_{\pm 0.1}$ $76.0$
DoPrompt
$69.5_{\pm 0.5}$ $59.1_{\pm 0.3}$ $79.9_{\pm 0.5}$ $82.7_{\pm 0.2}$ $72.8$
StochCA
$72.4_{\pm 0.8}$ $60.1_{\pm 0.5}$ $81.6_{\pm 0.2}$ $84.3_{\pm 0.2}$ $74.6$
@@
ERM
$67.6_{\pm 0.1}$ $23.9_{\pm 0.3}$ $54.2_{\pm 0.4}$ $16.7_{\pm 0.3}$ $69.1_{\pm 0.0}$ $54.9_{\pm 0.2}$ $47.8$
${}^{\dagger}$
$35.6_{\pm 0.5}$ $13.9_{\pm 0.2}$ $33.1_{\pm 0.8}$ $6.4_{\pm 0.2}$ $40.6_{\pm 0.4}$ $32.8_{\pm 0.3}$ $27.1$
${}^{\dagger}$
$45.3_{\pm 0.5}$ $13.5_{\pm 0.0}$ $45.8_{\pm 0.0}$ $7.7_{\pm 0.2}$ $49.6_{\pm 0.1}$ $43.8_{\pm 0.2}$ $34.3$
${}^{\dagger}$
$45.6_{\pm 0.1}$ $14.4_{\pm 0.2}$ $44.6_{\pm 0.6}$ $8.1_{\pm 0.0}$ $52.3_{\pm 0.3}$ $44.8_{\pm 0.5}$ $35.0$
${}^{\dagger}$
$59.3_{\pm 0.2}$ $20.8_{\pm 0.2}$ $46.6_{\pm 0.6}$ $11.3_{\pm 0.5}$ $63.7_{\pm 0.1}$ $47.0_{\pm 0.3}$ $41.3$
${}^{\dagger}$
$66.8_{\pm 0.2}$ $24.4_{\pm 0.1}$ $54.6_{\pm 0.2}$ $16.2_{\pm 0.2}$ $69.3_{\pm 0.1}$ $54.9_{\pm 0.3}$ $47.7$
${}^{\dagger}$
$65.8_{\pm 1.4}$ $23.8_{\pm 0.1}$ $54.7_{\pm 0.1}$ $17.2_{\pm 0.2}$ $69.0_{\pm 0.1}$ $54.9_{\pm 0.3}$ $47.6$
${}^{\dagger}$
$65.1_{\pm 0.1}$ $23.7_{\pm 0.1}$ $54.4_{\pm 0.1}$ $15.1_{\pm 0.3}$ $67.9_{\pm 0.0}$ $54.0_{\pm 0.1}$ $46.7$
${}^{\dagger}$
$67.0_{\pm 0.2}$ $23.8_{\pm 0.2}$ $54.0_{\pm 0.0}$ $15.9_{\pm 0.5}$ $69.0_{\pm 0.1}$ $54.6_{\pm 0.2}$ $47.4$
${}^{\dagger}$ $67.7_{\pm 0.2}$ $24.6_{\pm 0.1}$ $54.9_{\pm 0.1}$ $17.5_{\pm 0.2}$ $69.6_{\pm 0.3}$ $55.2_{\pm 0.5}$ $48.3$
DoPrompt
$66.6_{\pm 0.3}$ $23.4_{\pm 0.2}$ $54.5_{\pm 0.2}$ $16.2_{\pm 0.2}$ $69.2_{\pm 0.3}$ $54.5_{\pm 0.0}$ $47.4$
StochCA
$67.0_{\pm 0.3}$ $23.9_{\pm 0.9}$ $54.5_{\pm 0.3}$ $16.8_{\pm 0.3}$ $69.0_{\pm 0.2}$ $54.7_{\pm 0.4}$ $47.7$
Table 4: Detailed results on OfficeHome (left) and DomainNet (right). Performances marked with ${}^{\dagger}$ refer to the results reported in [29].

### v2

@@ 4.4 Experimental Results
ERM $86.9$ $79.2$ $73.0$ $\mathbb{47.8}$ $71.7$
IRM $83.9$ $79.5$ $72.3$ $27.1$ $65.7$
CDANN $82.3$ $79.1$ $72.3$ $34.3$ $67.0$
DANN $82.6$ $79.6$ $71.9$ $35.0$ $67.3$
GDRO $86.6$ $78.9$ $74.5$ $41.3$ $70.3$
CORAL $86.2$ $79.2$ $74.5$ $47.7$ $71.9$
MLDG $87.0$ $78.7$ $\mathbb{74.8}$ $47.6$ $72.0$
Mixup $\mathbb{87.4}$ $79.1$ $74.5$ $46.7$ $72.0$
MMD $86.9$ $79.8$ $74.5$ $47.4$ $72.2$
DoPrompt $88.1$ $80.4$ $76.0$ $48.3$ $73.2$
DoPrompt $86.7$ $78.9$ $72.8$ $47.4$ $71.5$
StochCA $\mathbb{87.4}$ $\mathbb{80.0}$ $74.6$ $47.7$ $\mathbb{72.4}$
@@
ERM $83.4$ $75.9$ $66.4$ $44.4$ $67.5$
MIRO $95.6$ $82.2$ $82.5$ $54.0$ $78.6$
StochCA $94.8$ $82.1$ $82.2$ $58.0$ $79.3$
MIRO+StochCA $\mathbb{96.4}$ $\mathbb{83.0}$ $\mathbb{85.1}$ $\mathbb{60.0}$ $\mathbb{81.1}$
Table 2: Domain generalization performance using ImageNet and CLIP ViT backbones. The best performance for each dataset is highlighted in bold. Performances marked with and refer to the results reported in [29] and [14], respectively.
@@
ERM $90.3_{\pm 0.8}$ $84.2_{\pm 0.5}$ $99.3_{\pm 0.2}$ $73.9_{\pm 3.2}$ $86.9$
IRM $89.1_{\pm 0.7}$ $79.1_{\pm 1.1}$ $98.3_{\pm 0.1}$ $68.2_{\pm 5.1}$ $83.9$
CDANN $87.3_{\pm 0.6}$ $80.1_{\pm 0.5}$ $98.5_{\pm 0.2}$ $63.1_{\pm 2.9}$ $82.3$
DANN $87.5_{\pm 0.4}$ $78.6_{\pm 0.9}$ $98.0_{\pm 0.2}$ $66.3_{\pm 1.7}$ $82.6$
GDRO $90.1_{\pm 0.6}$ $80.5_{\pm 1.0}$ $98.8_{\pm 0.1}$ $77.1_{\pm 1.5}$ $86.6$
CORAL $89.7_{\pm 0.3}$ $80.2_{\pm 1.0}$ $98.7_{\pm 0.2}$ $76.1_{\pm 1.6}$ $86.2$
MLDG $90.4_{\pm 0.5}$ $83.0_{\pm 0.7}$ $98.7_{\pm 0.0}$ $76.0_{\pm 0.2}$ $87.0$
Mixup $91.1_{\pm 0.1}$ $84.0_{\pm 0.3}$ $99.2_{\pm 0.2}$ $75.4_{\pm 0.5}$ $87.4$
MMD $89.9_{\pm 0.5}$ $80.8_{\pm 0.5}$ $98.7_{\pm 0.1}$ $77.1_{\pm 0.7}$ $86.6$
DoPrompt $91.1_{\pm 0.3}$ $83.0_{\pm 0.2}$ $99.6_{\pm 0.0}$ $78.7_{\pm 0.8}$ $88.1$
DoPrompt $89.9_{\pm 2.2}$ $84.1_{\pm 0.7}$ $98.9_{\pm 0.0}$ $74.1_{\pm 3.3}$ $86.7$
StochCA $90.8_{\pm 0.6}$ $82.9_{\pm 0.7}$ $98.9_{\pm 0.1}$ $77.1_{\pm 2.4}$ $87.4$
@@
ERM $96.2_{\pm 0.5}$ $64.4_{\pm 0.9}$ $77.3_{\pm 0.3}$ $79.0_{\pm 1.0}$ $79.2$
IRM $96.2_{\pm 0.5}$ $65.7_{\pm 0.8}$ $76.8_{\pm 0.4}$ $77.7_{\pm 0.9}$ $79.5$
CDANN $96.3_{\pm 0.2}$ $66.3_{\pm 0.5}$ $76.5_{\pm 0.6}$ $77.3_{\pm 0.5}$ $79.1$
DANN $97.7_{\pm 0.3}$ $66.5_{\pm 0.5}$ $76.0_{\pm 1.0}$ $78.0_{\pm 0.7}$ $79.6$
GDRO $97.1_{\pm 0.5}$ $65.2_{\pm 0.7}$ $75.1_{\pm 1.0}$ $78.3_{\pm 1.0}$ $78.9$
CORAL $96.2_{\pm 0.7}$ $66.0_{\pm 0.3}$ $76.8_{\pm 0.6}$ $78.0_{\pm 0.9}$ $79.2$
MLDG $95.7_{\pm 0.5}$ $65.9_{\pm 0.3}$ $76.3_{\pm 0.8}$ $78.3_{\pm 1.0}$ $78.9$
Mixup $97.1_{\pm 0.9}$ $65.4_{\pm 0.2}$ $76.1_{\pm 0.4}$ $78.0_{\pm 0.3}$ $79.1$
MMD $96.8_{\pm 0.1}$ $65.6_{\pm 1.0}$ $77.3_{\pm 0.6}$ $79.3_{\pm 1.5}$ $79.8$
DoPrompt $97.2_{\pm 0.1}$ $67.4_{\pm 0.5}$ $77.4_{\pm 1.0}$ $79.7_{\pm 0.5}$ $80.4$
DoPrompt $96.2_{\pm 0.6}$ $64.5_{\pm 0.8}$ $76.8_{\pm 1.5}$ $78.0_{\pm 1.1}$ $78.9$
StochCA $96.3_{\pm 0.3}$ $65.9_{\pm 0.8}$ $77.9_{\pm 0.6}$ $79.9_{\pm 0.7}$ $80.0$
Table 3: Detailed results on PACS (left) and VLCS (right). Performances marked with refer to the results reported in [29].
@@
ERM $69.8_{\pm 0.5}$ $59.4_{\pm 1.0}$ $79.6_{\pm 0.5}$ $83.1_{\pm 0.4}$ $73.0$
IRM $69.5_{\pm 0.1}$ $57.1_{\pm 0.3}$ $79.1_{\pm 0.1}$ $83.3_{\pm 0.1}$ $72.3$
CDANN $68.2_{\pm 0.4}$ $56.8_{\pm 0.4}$ $79.8_{\pm 0.4}$ $84.2_{\pm 0.0}$ $72.3$
DANN $68.7_{\pm 0.9}$ $55.8_{\pm 0.4}$ $79.2_{\pm 0.0}$ $83.9_{\pm 0.2}$ $71.9$
GDRO $71.2_{\pm 0.7}$ $60.9_{\pm 0.4}$ $81.1_{\pm 0.3}$ $84.7_{\pm 0.3}$ $74.5$
CORAL $72.3_{\pm 0.7}$ $60.9_{\pm 1.0}$ $80.4_{\pm 0.4}$ $84.3_{\pm 0.0}$ $74.5$
MLDG $72.3_{\pm 0.0}$ $60.0_{\pm 1.1}$ $82.2_{\pm 0.2}$ $84.7_{\pm 0.1}$ $74.8$
Mixup $72.7_{\pm 0.4}$ $59.7_{\pm 0.5}$ $80.9_{\pm 0.3}$ $84.6_{\pm 0.2}$ $74.5$
MMD $71.8_{\pm 0.2}$ $60.7_{\pm 1.0}$ $81.4_{\pm 0.3}$ $84.3_{\pm 0.1}$ $74.5$
DoPrompt $72.7_{\pm 0.5}$ $62.3_{\pm 0.6}$ $83.2_{\pm 0.1}$ $85.9_{\pm 0.1}$ $76.0$
DoPrompt $69.5_{\pm 0.5}$ $59.1_{\pm 0.3}$ $79.9_{\pm 0.5}$ $82.7_{\pm 0.2}$ $72.8$
StochCA $72.4_{\pm 0.8}$ $60.1_{\pm 0.5}$ $81.6_{\pm 0.2}$ $84.3_{\pm 0.2}$ $74.6$
@@
ERM $67.6_{\pm 0.1}$ $23.9_{\pm 0.3}$ $54.2_{\pm 0.4}$ $16.7_{\pm 0.3}$ $69.1_{\pm 0.0}$ $54.9_{\pm 0.2}$ $47.8$
IRM $35.6_{\pm 0.5}$ $13.9_{\pm 0.2}$ $33.1_{\pm 0.8}$ $6.4_{\pm 0.2}$ $40.6_{\pm 0.4}$ $32.8_{\pm 0.3}$ $27.1$
CDANN $45.3_{\pm 0.5}$ $13.5_{\pm 0.0}$ $45.8_{\pm 0.0}$ $7.7_{\pm 0.2}$ $49.6_{\pm 0.1}$ $43.8_{\pm 0.2}$ $34.3$
DANN $45.6_{\pm 0.1}$ $14.4_{\pm 0.2}$ $44.6_{\pm 0.6}$ $8.1_{\pm 0.0}$ $52.3_{\pm 0.3}$ $44.8_{\pm 0.5}$ $35.0$
GDRO $59.3_{\pm 0.2}$ $20.8_{\pm 0.2}$ $46.6_{\pm 0.6}$ $11.3_{\pm 0.5}$ $63.7_{\pm 0.1}$ $47.0_{\pm 0.3}$ $41.3$
CORAL $66.8_{\pm 0.2}$ $24.4_{\pm 0.1}$ $54.6_{\pm 0.2}$ $16.2_{\pm 0.2}$ $69.3_{\pm 0.1}$ $54.9_{\pm 0.3}$ $47.7$
MLDG $65.8_{\pm 1.4}$ $23.8_{\pm 0.1}$ $54.7_{\pm 0.1}$ $17.2_{\pm 0.2}$ $69.0_{\pm 0.1}$ $54.9_{\pm 0.3}$ $47.6$
Mixup $65.1_{\pm 0.1}$ $23.7_{\pm 0.1}$ $54.4_{\pm 0.1}$ $15.1_{\pm 0.3}$ $67.9_{\pm 0.0}$ $54.0_{\pm 0.1}$ $46.7$
MMD $67.0_{\pm 0.2}$ $23.8_{\pm 0.2}$ $54.0_{\pm 0.0}$ $15.9_{\pm 0.5}$ $69.0_{\pm 0.1}$ $54.6_{\pm 0.2}$ $47.4$
DoPrompt $67.7_{\pm 0.2}$ $24.6_{\pm 0.1}$ $54.9_{\pm 0.1}$ $17.5_{\pm 0.2}$ $69.6_{\pm 0.3}$ $55.2_{\pm 0.5}$ $48.3$
DoPrompt $66.6_{\pm 0.3}$ $23.4_{\pm 0.2}$ $54.5_{\pm 0.2}$ $16.2_{\pm 0.2}$ $69.2_{\pm 0.3}$ $54.5_{\pm 0.0}$ $47.4$
StochCA $67.0_{\pm 0.3}$ $23.9_{\pm 0.9}$ $54.5_{\pm 0.3}$ $16.8_{\pm 0.3}$ $69.0_{\pm 0.2}$ $54.7_{\pm 0.4}$ $47.7$
Table 4: Detailed results on OfficeHome (left) and DomainNet (right). Performances marked with refer to the results reported in [29].